OpenStudy (abbycross167):

can someone please help me answer this question? I will become a fan and will give a medal! Question~ Solve the systems using the substitution method. y = (2/3)x + 1 y + x = 6

OpenStudy (anonymous):

so all you have to figure out is what is the value of x which will show the value of y

OpenStudy (abbycross167):

ok, could you shoe me how to do that? Please?

OpenStudy (anonymous):

Just keep trying equations that add up to 6. Kind of like 3+3

OpenStudy (abbycross167):

I don't really understand what you mean...

OpenStudy (anonymous):

yeah the answer is 3 because \[(2\div3)\times3+1 = 3\]

OpenStudy (anonymous):

Basicly replace x with 3 Then the answer shows that y is three And when you add those together it's 6 ^-^

OpenStudy (anonymous):

so the x and y value is 3

OpenStudy (abbycross167):

Oh ok, Can I show you the example from my lesson and we can try to figure it out using those steps? because I have to show my steps...

OpenStudy (anonymous):

Yeah

OpenStudy (abbycross167):

Method 2 Substitution Example 2: Solve the following system of linear equations by graphing. 2x - 3y = -2 4x + y = 8 Solution: S1 Solve one equation for either variable. Choose the second equation since y is by itself: y = 8-4x S2 In the other equation, replace the variable from S1 with what it equaled. 2x - 3(8-4x) = -2 S3 Solve the equation in S2 for the other variable. 14x = -2 + 24 -- x = 22/14 = 11/7 S4 Substitute the value found in S3 into equation in S1. y = 8 - 4(11/7) S5 Solve the equation in S4. y = 8 – (44/7) -- y = 12/7 S6 Write your answer as an ordered pair (x, y). (11/7, 12/7)

OpenStudy (abbycross167):

thats the example from the lesson

OpenStudy (anonymous):

yeah just use those steps srry i gtg brb

OpenStudy (abbycross167):

umm ok